# Fixed-domain curve counts for blow-ups of projective space

### v3

@@ 4.2. The quantum Euler class of $X$
where in the second equality we have used Lemma 57, in the third equality we have written $\mathsf{H}=\mathsf{H}-\mathsf{E}+\mathsf{E}$, and in the fourth equality we have used the first relation in (20) and the fact that, by Lemma 57, we have
@@ 4.3. Proof of Theorem 14
and the proof of Theorem 14 is reduced to the following.
Lemma 59.
@@
unless $\ell-d-m-1\geq k$, which, using (22), is equivalent to $d\geq(2r-1)k$. In particular, we get 00 unless $\ell\geq r+2$.
Proof of Lemma 59 .
@@
In particular, the coefficient of $\mathsf{P}q^{d\mathsf{H}^{\vee}+(k+1)\mathsf{E}^{\vee}}$ is 00. The claimed equality then follows from Remark 60.
@@
This concludes the proof of Theorem 14.
Remark 61.
The condition $n-d\geq 1$ in Theorem 14 (i.e., the condition $\ell-m-d\geq 1$ in Lemma 59 ) is necessary. For example, when $r=2$ and $n=g=d=k=1$, we have
@@
while the right-hand side of Theorem 14 is equal to $1$.
Remark 62.
When $k=0$, in Lemma 59 we have

### v4

@@ 4.2. The quantum Euler class of $X$
where in the second equality we have used Lemma 4.3, in the third equality we have written $\mathsf{H}=\mathsf{H}-\mathsf{E}+\mathsf{E}$, and in the fourth equality we have used the first relation in (24) and the fact that, by Lemma 4.3, we have
@@
and the proof of Theorem 1.14 is reduced to the following.
Lemma 4.5.
@@
unless $\ell-d-m-1\geq k$, which, using (26), is equivalent to $d\geq(2r-1)k$. In particular, we get 0 unless $\ell\geq r+2$.
Proof of Lemma 4.5 .
@@
In particular, the coefficient of $\mathsf{P}q^{d\mathsf{H}^{\vee}+(k+1)\mathsf{E}^{\vee}}$ is 0. The claimed equality then follows from Remark 4.6.
@@
This concludes the proof of Theorem 1.14.
Remark 4.7.
The condition $n-d\geq 1$ in Theorem 1.14 (i.e., the condition $\ell-m-d\geq 1$ in Lemma 4.5 ) is necessary. For example, when $r=2$ and $n=g=d=k=1$, we have
@@
while the right-hand side of Theorem 1.14 is equal to $1$.
Remark 4.8.
When $k=0$, in Lemma 4.5 we have